Kyky232:

In parallelogram EFGH, the measure of angle F is (3x − 10)° and the measure of angle G is (5x + 22)°. What is the measure of angle G? 42° 53° 117° 127°

darkknight:

Since these are 2 angles that are like this  They will both add up to 180 degrees so then 3x-10) + (5x+22) = 180 8x + 12 = 180 8x = 192 x = ? And then plug x back into 5x+22 and find the angle of G

Kyky232:

So I did something wrong... I got 142 and thats not an option

darkknight:

oops sorry, im a big brain

darkknight:

3x-10) + (5x+22) = 180 8x + 12 = 180 We subtract 12 from both sides, not add..... 8x = 168 x = ? And then plug x back into 5x+22 and find the angle of G

Kyky232:

So its 127?

darkknight:

Yep, have a good day :)

Kyky232:

You too!